# Magnetic levitation by rotation described by a new type of Levitron

### v1

@@ V.3 The condition on the maximum allowed value of lateral displacement
In our experiments, the magnet coupled to the rotor consisted of a cylindrical magnet with height and diameter $d=h=5mm$, assuming that the magnetization of the cylinder is uniform, we can represent the magnetic field at the center of the magnet can be expressed as:
@@
In FIG.(6) we show how the radius $m/\mu_{f}$ varies as a function of $\delta_{R}$. The region shaded in purple delimits the results obtained from direct measurements of the magnetic field at a distance $r_{f}=10mm$, with the minimum and maximum values $B_{min}=2.1mT$ and $B_{max}=6.5mT$, respectively. In this figure, the solid red line corresponds to the direct determination of $m/\mu_{f}=0.01158Kg/A.m^{2}$. From this plot, we find $0.35mm\gtrsim\delta^{c}_{R}>0.1mm$ which agrees with the results found in subsection A.
@@
where $a(r_{f})=h^{3}/r^{3}_{f}$ and $b(r_{f})=4/r_{f}$. Measurements of $B(r_{f})$ and $r_{f}$ allow the identification of $m/\mu_{f}(\delta_{R})$, for a given specific rotor configuration defined by $\delta_{R}$. Assuming limiting case where $R_{f}=R_{f(max)}\to 1$ , by extrapolating from the 3D contour plot provided at the end of Appendix A we can then identify for $B_{min}(10mm)=2.1mT$ and $B_{max}(3mm)=0.125T$, where $3mm\lesssim r_{f}\lesssim 10mm$ , that
@@
VI Conclusions
In this work, considering the results reported in subsections IV-A to IV-C, we present evidence suggesting that the novel type of magnetic levitation described in Refs.[1][2], can in fact be interpreted as a magnetic levitation arising from a new variant of the Levitron, namely the ”dynamical Levitron”.
The main difference between these two systems, as previously discussed lies in the fact that for the dynamical levitron the radius $m_{/}\mu_{f}$ depends on the lateral displacement $\delta_{R}$. Consequently, the adjustment required for magnetic trapping is not determined by the rotation speed of the floater, but rather by the appropriate choice of $\delta_{R}$, allowing for trapping in which the float can be at rest, as shown in the videos available at[7] .
By considering an effective model, we found that the observed levitation effect can be explained within the context of the discussion presented in Refs.[3][4], where for the potential energy to be a minimum at the trapping point, we verify the existence of a condition under the $\delta_{R}$ parameter. As a consequence, as shown throughout sections IV and V, the occurrence of the magnetic trapping effect of the floater is primarily due to the condition that the lateral displacement $\delta_{R}$, being observed the trapping when $\delta_{R}\gtrsim\delta^{c}_{R}$.
@@ VI Conclusions
In some situations involving magnetic trapping we also identified distinct vibrational modes. However, after noticing that the rotor and floater speeds became strongly correlated at the moment of floater trapping, we verified that instabilities and vibrations arising from the rotor ended up being transmitted at the moment of coupling to the floater, which were then removed with the appropriate mechanical isolation of the motor used in our experimental setup. The data presented in this work, including videos lasting more than 2 hours, which demonstrate the magnetic stability presented in the experiments carried out, are available at Ref.[7].
@@ Acknowledgements.
Appendix A Tables

### v3

@@ 5.5 The condition on the maximum allowed value of lateral displacement
In our experiments, the magnet coupled to the rotor consisted of a cylindrical magnet with height and diameter $d=h=5mm$, assuming that the magnetization of the cylinder is uniform, the magnetic field at the center of the magnet can be expressed as:
@@
In FIG.(6) we show how the radius $m/\mu_{f}$ varies as a function of $\delta_{R}$. The region shaded in purple delimits the results obtained from direct measurements of the magnetic field at a distance $r_{f}=10mm$, with the minimum and maximum values $B_{min}=(2.1\pm 0.1)mT$ and $B_{max}=(6.5\pm 0.3)mT$, respectively. In this figure, the solid red line corresponds to the direct determination of $m/\mu_{f}=0.01158Kg/A.m^{2}$. From this plot, we find $0.35mm\gtrsim\delta^{c}_{R}>0.1mm$ which agrees with the results found in subsection 5.1.
@@
where $a(r_{f})=h^{3}/r^{3}_{f}$ and $b(r_{f})=4/r_{f}$. Measurements of $B(r_{f})$ and $r_{f}$ allow the identification of $m/\mu_{f}(\delta_{R})$, for a given specific rotor configuration defined by $\delta_{R}$. Assuming limiting case where $R_{f}=R_{f(max)}\to 1$ , by extrapolating from the 3D contour plot provided at FIG.(7) we can then identify for $B_{min}(10mm)=(2.1\pm 0.1)mT$ and $B_{max}(3mm)=(0.125\pm 0.006)T$, where $3mm\lesssim r_{f}\lesssim 10mm$ , that
@@
6 Conclusions
In Refs.[1][2], the authors established the fundamental physical principles underlying the magnetic levitation observed in conventional Levitron systems. These works demonstrate that stable levitation necessitates a precise tuning of the ratio $m_{/}\mu_{f}$, which ensures the system achieves a potential minimum at the trapping point.
Building upon this established framework and considering the results from subsections V-1 to V-5, this paper extends these principles to a ”dynamical Levitron” [4][6]. The main difference between these two systems, as previously discussed, is that in the dynamical Levitron, the ratio $m_{/}\mu_{f}$ becomes dependent on the lateral displacement $\delta_{R}$. Consequently, the adjustment required for magnetic trapping is not governed by the rotation speed of the floater, but rather by the appropriate choice of $\delta_{R}$. This distinct characteristic enables trapping configurations where the floater can remain at rest, as shown in the videos available at[7].
By considering an effective model, we found that the observed levitation effect can be explained within the context of the discussion presented in[1][2], where for the potential energy to be a minimum at the trapping point, we verify the existence of a condition under the $\delta_{R}$ parameter. As a consequence, as shown throughout sections IV and V, the occurrence of the magnetic trapping effect of the floater is primarily due to the condition that the lateral displacement $\delta_{R}>0$, being observed the trapping when $\delta_{R}\gtrsim\delta^{c}_{R}$.
@@ 6 Conclusions
In some situations involving magnetic trapping we also identified distinct vibrational modes. However, after noticing that the rotor and floater speeds became strongly correlated at the moment of floater trapping, we verified that instabilities and vibrations arising from the rotor ended up being transmitted at the moment of coupling to the floater, which were then removed with the appropriate mechanical isolation of the motor used in our experimental setup.
@@ Acknowledgments
Data availability statement
All data that support the findings of this study are included within the article (and any supplementary material in [7]). It is important to highlight that [7] refers to a repository containing all the video recordings of the experiments conducted in this study.
Appendix A Experimental Setup
Figure 8: The experimental setup: Motor used in the construction of the rotor and floats used.
We have performed an experimental setup as follows. As depicted in Figs.(A1-a),(A1-b) , a Nd-Fe-B cylindrical magnet with diameter of $d=5mm$ and height $h=5mm$, whose nominal remanence is (1.171.22)T(1.17–1.22)T, was firmly attached to the motor shaft using super bonder glue.
Coaxially with this magnet, a disk of radius $r=15mm$ was fixed to enable real-time measurement of the motor’s rotational speed by a digital tachometer’s laser sensor, as shown in Fig.(A2-a), that is used to determine the motor rotational speed in real time.
The motor identified in Fig.(A1), consists of a micro motor DC whose operating specifications correspond to $V_{max}=24V$, $i_{max}=1.21A$ and $\omega_{max}=21,600rpm$. The floaters employed in the experiments, documented in the videos available in [7] and illustrated in Fig. (A1-c) [^2] Note that in the preparation of this photo, the magnets were placed under a small metal plate inserted below a sheet of white paper., were marked to facilitate the determination of $\omega_{f}$ using the motion-tracking software Tracker.
To control the motor’s speed, a stabilized power source, depicted in Fig. A2-b, was employed. This source allows for precise adjustment of voltage (V) and current (I), thereby controlling its rotational speed $($$\omega$). As detailed in the tables provided in Appendix B, the motor rotational speeds required for the magnetic trapping of the float typically range from $\omega_{f}(10,400-13,700)rpm$. The intrinsic alternating magnetic field of the motor at a typical distance of the floater magnet is assumed to be negligible.
Figure 9: The experimental setup: Devices used in the characterization of measurements.
Figure 10: The experimental setup: Lateral displacements $\delta_{R}=1mm$ (a) and $\delta_{R}=1.75mm$ (b).
The distinct configurations for lateral displacement $\delta_{R}$ that we consider in this work were incorporated in different arrangements similar to those shown in Fig.(A1). For illustration purposes, in Fig.(A3) we present the configuration in which $\delta_{R}=1mm$ , Fig.(A3-a) and $\delta_{R}=1.75mm$ in Fig.(A3-b).
Figures A4-a and A4-b illustrate the methodology employed to obtain the tabulated results in Appendix B. During experiments, the dynamical behavior of the floater magnet was recorded and these video recordings were subsequently post-processed using the motion-tracking software Tracker, in order to determine the position and orientation of the floater as a function of time.
Figure 11: The experimental setup: High-resolution camera footage at 920 fps of the magnet levitating.
As we mentioned, the determination of the rotor(R) rotation speed, $\omega_{R}$, is obtained from the direct reading provided by the digital tachometer. The specific instrument used for these measurements is a Minipa Mdt-2244c digital photo tachometer. This device features a rotational speed range from $\omega_{min}=10rpm$ to $\omega_{max}=99,999rpm$, and the minimum and maximum distance range between the sensor and the rotating surface corresponds to (50200)mm(50-200)$m$. The accuracy in determining the rotational speed, for a reading falling within the range (50200)mm(50-200)$m$ , corresponds to $\Delta\omega=\pm 0.04\%$.
In turn, the floater’s rotational speed, $\omega_{f}$, was determined through the analysis of high-resolution slow-motion videos using the Tracker software. These recordings were captured at 960 frames per second (fps) with a Samsung Galaxy S23 mobile phone. The error in Tracker analysis varies significantly by application, but common sources of error include video quality (resolution, blur), calibration accuracy, object visibility, and the precision of manual or automatic point marking.
There is no single ”error” value for Tracker; instead, it depends on specific experimental conditions. Manually marking the center of an object in each frame in videos is a source of uncertainty, that can estimate as $O(1)\%$, and can be assumed in the our estimates for $\omega_{f}$.
Finally, the digital Gaussmeter WT10A represented in Fig.(A2-c), was used to characterize the magnetic field measurements reported in section V-5, the magnetic field measurement range for this device corresponds to (02000)mT(0-2000)mT, presenting an accuracy of $\Delta B=\pm 5\%$. Considering the resolution of this device, the magnetic field measurements reported in subsection V-5 correspond to $B_{min}=(2.1\pm 0.1)mT$ and $B_{max}=(6.5\pm 0.3)mT$.
Appendix B Tables